# Everything Everywhere All At Once: A First Principles Numerical Demonstration of Emergent Decoherent Histories

### v2

@@ III.2 Results
IV Consequences
For coarse and slow observables in isolated non-integrable many-body systems, the finite size scaling observed above shows an exponential suppression of quantum effects and that this suppression is typical, thus strongly suggesting that quantum interference effects vanish for macroscopic systems. Based on an arguably minimal set of assumptions this provides an explanation why the MWI is compatible with a classically perceived reality, provided we focus on the perception of slow and coarse observables (as we humans do). Experimentally, for any fixed measurement error the figures above can be used to interpolate at which point the coarse dynamics of an isolated quantum system becomes effectively indistinguishable from a classical stochastic process.
Cleary, there are many details to be filled in (see Sec. V), but in this section we take the perspective that the above theory is correct and explore its consequences. In particular, we consider from that perspective the set selection problem, the initial state of the Universe, the foundations of quantum mechanics and the arrow of time.
IV.1 Set selection problem
The set selection problem has been used to criticize the histories formalism by pointing out that the DHC is incomplete Paz and Zurek (1993); Dowker and Kent (1995, 1996); Riedel et al. (2016), see also Refs. Zurek (2003); Schlosshauer (2019); Zurek (2022). The basic reasoning behind the criticism is the following. Consider for fixed $L$ and $M$ the manifold H of all possible projectors that can be used to define histories x (of very different physical meaning). The dimension of this manifold scales like $D^{2L}$, whereas the number of constraints imposed by the DHC in Eq. (3) scales like $M^{2L}$. Thus, the submanifold DH of decoherent histories has a dimension that scales like $D^{2L}-M^{2L}$, which (even though of measure zero with respect to H) is enormous for $D\gg M$. Hence, which of the many physically very distinct histories should one use? [^6] Note the close similarity to the preferred basis problem.
One response to this problem is to claim that this is actually not a problem Griffiths (2019), but we would argue that there is no set selection problem if one focuses on slow and coarse observables of many-body systems (that is, situations relevant to human perception), and any remaining ambiguity is actually important. While this has been remarked in Ref. Griffiths (2019), a systematic explanation backed up by unambiguous evidence has not been given.
The crucial insight is to demand a certain stability or robustness of decoherence with respect to the initial state $|\psi(t_{0})\rangle$ and times $t_{k}$. To repeat: above we found approximate decoherence (without exception) for slow and coarse observables for a large set of initial states, times and Hamiltonians. In contrast, it seems likely that the vast majority of decoherent histories in DH is not robust in this sense. Instead, it likely depends sensitively on, e.g., $|\psi(t_{0})\rangle$, as sketched in Fig. 11.
@@ IV.1 Set selection problem
One might object that those examples are nevertheless legitimate, but there are at least two fundamental practical obstacles. First, it is unclear how humans should get access to a projector of the form $|\psi(t_{k}){\rangle\langle}\psi(t_{k})|$ using earthly instruments. Second, proper relativistic considerations teach us that at any given time $t$ we can only access part of the universal wave function $\rho_{\text{accessible}}(t)=\mbox{tr}_{\text{inaccesible}}\{|\psi(t){\rangle\langle}\psi(t)|\}$. Thus, even if we assume we find (by some magic instrument) a strange decoherent history depending on $\rho_{\text{accessible}}(t)$, decoherence would likely be destroyed in the next second once a new photon from an hitherto unobserved star reaches us.
@@
Second, we repeat that the proposed solution is restricted to slow and coarse observables of many-body systems that approximately decohere. It does not solve the set selection problem if one demands exact decoherence. Also other fundamental problems within the histories formalism such as the one debated in Refs. Kent (1997); Griffiths and Hartle (1998); Kent (1998) are not directly influenced by our argument.
IV.2 Initial state of the Universe
Our current understanding of the Universe is based on two separate concepts: dynamical laws (Newton, Maxwell, Schrödinger, Einstein, etc.) on the one hand, and boundary conditions—often in form of initial conditions—on the other hand. Therefore, much effort is devoted to understand the “state of the Universe”, often at some “initial” time where things were particularly simple.
This is true quantum mechanically and classically. For instance, the arrow of time and the second law suggest for a classical Universe that it must have started in an extremely exceptional and fine tuned low entropy state Feynman et al. (1963); Penrose (1989); Lebowitz (1993); Price (1996); Albert (2000), together with controversies as to whether this fine tuning requires further explanation or not Price (2004); Callender (2004). Moreover, the smoothness of the cosmic microwave background, as explained by the theory of inflation, suggests that quantum features have played a decisive role in the early Universe. Consequently, the wave function of the Universe plays an important role for various problems Hartle and Hawking (1983); Hawking (1984); Tipler (1986a); Mukhanov (2005); Carr (2007); Lehners (2023).
If the exponential suppression of quantum effects is typical, as predicted by the theory presented here, this suggests that the quest for the wave function of the Universe is more subtle than one might hope for, for two reasons. First, according to the MWI (together with an emergent classical structure) all we can infer is that the wave function of the Universe has some unknown but non-zero component compatible with our branch or history, an evident point made already before, e.g., in Ref. Tipler (1986b). Second, we also found that the emergence of decoherence is typical, i.e., almost all possible wave functions give rise to the same macroscopic Multiverse from a statistical point of view, i.e., up to random short fluctuations and unlikely statistical outliers. To see this, recall that our results were obtained for an initial equilibrium state and that one can mathematically rigorously show, using the concept of measure concentration (typicality), that almost all states are in equilibrium Farquhar and Landsberg (1957); Bocchieri and Loinger (1958); Gemmer et al. (2004); Popescu et al. (2006). Moreover, even when starting from an unlikely nonequilibrium macrostate, this state still lives in an enormously large subspace (according to Penrose’s estimate Penrose (1989) the entropy at the time of the big bang was around $10^{88}k_{B}$, leaving us still with on the order of ${\cal{O}}(10^{10^{88}})$ many microstates) and will typically give rise to decoherence.
Our results therefore imply that the emergence of decoherence requires no fine tuning. [^9] Note that we only refer to the fine tuning problem of the initial state. The fine tuning problem of the fundamental constants is not solved within the present considerations (see Ref. Carr (2007) for further discussion). Moreover, even a Universe in equilibrium gives rise to a Multiverse with many interesting nonequilibrium universes, even though their probability can be very small (we discuss this further in Sec. IV.4). Thus, at least without invoking further arguments, observing the fact that we can reason about our (local) universe classically implies almost nothing for the global wave function of the Universe.
IV.3 Foundations of quantum mechanics
We have just found that, as far as the emergence of decoherence is concerned, the MWI applied to slow and coarse observables gives rise to classical universes for almost all initial states. These initial states might be at equilibrium, very far from equilibrium (as the current big bang model suggests), or anything in between. Moreover, Sec. IV.1 argued that the histories framework, perhaps combined with other arguments such as quantum Darwinism, provides enough structure to reconcile the quantum Multiverse with our experience of living in one universe. It is thus scientifically questionable to reject the MWI only because of its seemingly “absurd” consequences. The MWI removes at least part of the fine tuning problem associated to the initial state of the Universe and it provides a basis for anthropic reasoning because there is something to select from (for a broad discussion see Ref. Carr (2007)).
These features of the MWI are not shared in any obvious way by interpretations of quantum mechanics that postulate a real objective (and not only apparent) collapse of the wave function due to measurements or hidden variables. Since these interpretations imply the existence of one single Universe, it seems likely that they face difficulties to explain the origin of our Universe without invoking a lot of lucky chance. This observation is interesting from a historical perspective because Ockham’s razor was a common objection against the MWI Shimony (1963): Why should one introduce a very large number of worlds that nobody can see? But not only can this argument be reversed—why and when should one introduce a large number of unnecessary collapses? Vaidman (2021)—but the MWI might even provide a simpler basis for cosmological considerations, see also Refs. Hartle and Hawking (1983); Hawking (1984); Tipler (1986a); Mukhanov (2005); Carr (2007); Lehners (2023).
IV.4 Arrow(s) of time
The Schrödinger equation obeys time reversal symmetry, and the initial equilibrium state used in the above simulation also does not introduce any time asymmetry. It is therefore an important consistency check Price (1996) that one finds (statistically) as many histories with a forward arrow of time as a backward one, as demonstrated in Fig. 6. Thus, arrows of time can be an emergent concept in a time reversal symmetric Universe (in varying contexts this point was also made in Refs. Carroll and Chen (2004); Barbour et al. (2014); Deutsch and Aguirre (2022)). Moreover, we remark that the DF can be as well defined with respect to the time-reversed evolution describing a branching of the wave function backwards in time (which would be natural for quantum retrodiction Barnett (2014)). Finally, there is actually also no preferred “initial” time $t_{0}$ in our model: we would observe the same phenomenology for any choice of $t_{0}$, a fact that is again related to our initial equilibrium state and the typicality of decoherence, see also Fig. 3 (b). Some special time $t_{0}$ could have been justified if one knows that the wave function of the Universe is at that time special too (i.e., out of equilibrium), but even then this would be a pure convention helpful for some calculations but not all.
In summary, our model demonstrates a time reversal symmetric system where arrows of time form on individual branches of the MWI. Unfortunately, owing to the restriction to only three macrostates $x\in\{-,0,+\}$, these arrows of time are minimal and do not suffice to explain our real world. While this problem can be overcome by considering more macrostates, branches further away from equilibrium will have less probability. This is a quantum version of the classical Boltzmann brain paradox (see also Ref. Albrecht and Sorbo (2004) for a discussion in context of inflation) that is often invoked to argue against the possibility that our universe is a mere fluctuation Carroll (2020), but a different reading is possible Müller (2020). Our model is still too simple to contribute to this debate, but it rigorously demonstrates at a quantitative level that classical universes with different symmetrically distributed arrows of time, even if they are mini-arrows, can emerge from equilibrium. This point is further discussed in Ref. Albrecht et al. (2022).
To better incorporate time reversal symmetry, one could modify the DF Isham (1994); Isham and Linden (1997); Griffiths (2002) or use other time symmetric formulations Aharonov et al. (1964); Vaidman (2010). Such a time symmetric view would then contradict common presentations of the MWI, where an exponentially increasing branching of the wave function occurs in one time direction (usually aligned with the authors’ own arrow of time). Indeed, some works went even further and identified this one-directional branching of the wave function with the increase in thermodynamic entropy Deutsch (2010); Aaronson (2013). In our view, there is nothing that warrants this conclusion, and it does not explain the apparent arrow of time.
@@ V Concluding Perspectives
We presented a direct evaluation of the DF for multiple (up to five) time steps from first principles for a non-trivial example. We rigorously defined quantum interference effects and extracted a scaling law of the form $D^{-\alpha}$ by varying the Hilbert space dimension $D$ over four orders of magnitude. This provides a firm starting point to quantitatively discuss the DHC.
The resulting picture is that coarse and slow observables of non-integrable many-body systems give rise to a robust and stable form of approximate decoherence; see also Refs. Van Kampen (1954); Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Albrecht et al. (2022); Strasberg et al. (2023); Strasberg (2023) for supporting evidence. We have further argued that this robustness and stability could solve the preferred basis problem of the MWI or the set selection problem of the histories formalism, respectively.
@@
In fact, chaos has been even realized as detrimental for the emergence of classicality within the Wheeler-DeWitt equation due to the breakdown of the WKB approximaton and EID was invoked to remedy for that Calzetta and Gonzalez (1995); Cornish and Shellard (1998); Calzetta (2012). Within our non-relativistic quantum mechanical model we can not make any direct contribution to this question, but one view suggested by this work is that emergence of classicality is best viewed as a synergy of different mechanisms instead of a single all-ruling idea. Indeed, it has been recently pointed out that the emergence of classicality in the early Universe is still an unsolved puzzle Berjon et al. (2021), and the present perspective might be able to add a piece to it.
Besides providing a firm starting point to quantitatively discuss the DHC, our work also shows that it is possible to discuss a variety of further aspects quantitatively, for instance, the distribution of arrows of time, the number of Boltzmann brains or typicality aspects of the initial wave function of the Universe. More broadly speaking, our results improve the understanding of multi-time quantum processes, many-body physics and pure state statistical mechanics.
@@
Also in these contexts many open questions remain. For instance, it would be desirable to get a better analytical understanding of the DHC. The results reported in Refs. Strasberg et al. (2023); Strasberg (2023) were restricted to three-time correlation functions and generalizing them to higher orders likely requires novel techniques (however, some progress for arbitrary $L$-time correlation functions in a different context was reported in Refs. Figueroa-Romero et al. (2019, 2021); Dowling et al. (2023a, b)). Further insights about the time scales at which decoherence arises are also desirable, and it is in particular also worth to ask about “recoherence times”: due to Poincaré recurrences the emergence of decoherence in a finite dimensional quantum system can not be a persistent, eternal phenomenon. In particular, it would be intriguing to find out how the recoherence time scales with the number of time steps $L$ or, more generally, with the net information aquired along a history. Whether non-integrability is really necessary for the robust emergence of decoherence from first principles (in particular without using ensemble averages) is another largely open question.
@@
Finally, our model calculations presently suggest that for any coarse and slow observable of a non-integrable system every possible classical history is realized. While this suggests the realization of a wide array of histories, it is not necessarily true that “everything that can happen will happen” (as sometimes portrayed in both scientific and popular accounts of the MWI). For instance, extremely unlikely (sequences of) outcomes might be restricted to very small subspaces because they are highly atypical, but our results indicate that large subspaces are needed for the emergence of classicality, see also Ref. Strasberg and Schindler (2023). Moreover, whether all the perceived randomness in our world stems from quantum effects Albrecht and Phillips (2014) or not Del Santo and Gisin (2019) remains a fascinating question too.
@@ Acknowledgements
We are grateful to Anthony Aguirre, Josh Deutsch and Jonathan Halliwell for discussions. PS is financially supported by “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/PR21/11840014) and acknowledges further support from the European Commission QuantERA grant ExTRaQT (Spanish MICINN project PCI2022-132965), by the Spanish MINECO (project PID2019-107609GB-I00) with the support of FEDER funds, the Generalitat de Catalunya (project 2017-SGR-1127) and by the Spanish MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1). JS acknowledges support by MICIIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by the Generalitat de Catalunya.

### v4

@@ III.2 Results
IV Set Selection Problem
The set selection problem has been used to criticize the histories formalism by pointing out that the DHC is incomplete Paz and Zurek (1993); Dowker and Kent (1995, 1996); Kent (1998a); Zurek (2003); Riedel et al. (2016); Schlosshauer (2019); Zurek (2022). The basic reasoning behind the criticism is the following. Consider for fixed $L$ and $M$ the manifold H of all possible projectors that can be used to define histories x (of very different physical meaning). The dimension of this manifold scales like $D^{2L}$, whereas the number of constraints imposed by the DHC in Eq. (3) scales like $M^{2L}$. Thus, the submanifold DH of decoherent histories has a dimension that scales like $D^{2L}-M^{2L}$, which (even though of measure zero with respect to H) is enormous for $D\gg M$. Hence, which of the many physically very distinct histories should one use? Note the close similarity to the preferred basis problem.
We here discuss the set selection problem based on the robust numerical findings from above (for further discussion see also Ref. Griffiths (2019)). In particular, and contrary to previous claims, we suggest that there simply is no set selection problem if one focuses on slow and coarse observables of many-body systems—that is: situations relevant to human perception—and the remaining freedom in the choice of observable is actually important for the ability of different observers to agree.
The key to our argument is to demand a certain stability or robustness of decoherence with respect to the initial state $|\psi(t_{0})\rangle$ and times $t_{k}$ and to focus on approximate instead of exact decoherence, which is sufficient for all practical purposes. To recall our results above: we found approximate decoherence without exception for slow and coarse observables for a large set of initial states, times and Hamiltonians. In contrast, it seems likely that the vast majority of decoherent histories in DH is not robust in this sense. Instead, it likely depends sensitively on, e.g., $|\psi(t_{0})\rangle$, as sketched in Fig. 12.
@@ IV Set Selection Problem
One might object that those examples are nevertheless legitimate, but there are fundamental practical obstacles. First, it is unclear how humans should get access to a projector of the form $|\psi(t_{k}){\rangle\langle}\psi(t_{k})|$ using earthly instruments. Second, if the MWI is correct, then knowledge of the global $|\psi(t_{k})\rangle$ is impossible since observers are limited to their own branch $|\psi(\textbf{x})\rangle$. Finally, proper relativistic considerations teach us that at any given time $t$ we can only access part of the universal wave function $\rho_{\text{accessible}}(t)=\mbox{tr}_{\text{inaccesible}}\{|\psi(t){\rangle\langle}\psi(t)|\}$. Thus, even if we assume we find (by some magic instrument) a strange decoherent history depending on $\rho_{\text{accessible}}(t)$, decoherence would likely be destroyed in the next second once a new photon from an hitherto unobserved star reaches us.
@@
Second, we repeat that the proposed solution is restricted to slow and coarse observables of many-body systems that approximately decohere. It does not solve the set selection problem if one demands exact decoherence. Also other fundamental problems within the histories formalism such as the one debated in Refs. Kent (1997); Griffiths and Hartle (1998); Kent (1998b) are not directly influenced by our argument.
@@ V Concluding Perspectives
V.1 Summary
We presented a direct evaluation of the DF for multiple (up to five) time steps from first principles for a non-trivial example. We rigorously defined quantum interference effects and extracted a scaling law of the form $D^{-\alpha}$ by varying the Hilbert space dimension $D$ over four orders of magnitude. This was checked for a wide variety of situations and it provides a firm starting point to quantitatively discuss the DHC.
The resulting picture is that coarse and slow observables of non-integrable many-body systems give rise to a robust and stable form of approximate decoherence. While our calculations were restricted to a particular model, the success and versatility of random matrix theory suggests that our results are more widely applicable; see also Refs. Van Kampen (1954); Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Albrecht et al. (2022); Strasberg et al. (2023); Strasberg (2023) for supporting evidence.
Based on this evidence, we suggested that the preferred basis problem of the MWI or the set selection problem of the histories formalism, respectively, is solvable if one is interested in observables relevant to us humans (which are slow and coarse and give rise to a robust and stable form of decoherence) and if one focuses on approximate instead of exact decoherence. Indeed, our results indicate that for such observables the emergence of decoherence is universal and happens for almost all pure states: no product state assumption, no low entropy initial state and no ensemble average was necessary in our approach. Moreover, it seems reasonable to expect (but difficult to prove) that possible ambiguities in the choice of branches/histories become irrelevant as different slow and coarse obervables approximately commute for large quantum systems.
@@ V.1 Summary
Finally, we also explicitly saw that an equilibrated Multiverse gives rise to branches with locally well-defined entropic arrows of time while overall the Multiverse remains statistically time-symmetric. While our model could be claimed to be unrealistic because of the Boltzmann brain paradox Albrecht and Sorbo (2004); Carroll (2020); Müller (2020), it nevertheless explicitly illustrates two crucial features. First, the branching structure of the Multiverse is pure convention and does not explain our arrow of time or meaningful thermodynamic entropy, contrary to suggestions made in Refs. Deutsch (2010); Aaronson (2013). Instead, from a fundamental point of view it would be preferable to start from a time-symmetric description of the MWI or histories formalism Aharonov et al. (1964); Isham (1994); Isham and Linden (1997); Griffiths (2002); Vaidman (2010). Second, opposite arrows of time can peacefully coexist within the same quantum Multiverse in contrast to opposite arrows of time in space-like separated regions of a classical Multiverse Schulman (1999), which are unstable with respect to the slightest perturbation Kupervasser et al. (2012). For further details about the emergence of classicality in time-symmetric situations see Ref. Albrecht et al. (2022). Moreover, that arrows of time can be an emergent concept in a time reversal symmetric Universe was also noted in different contexts, e.g., in Refs. Schulman (1999); Carroll and Chen (2004); Barbour et al. (2014); Deutsch and Aguirre (2022).
V.2 Outlook
Our work could stimulate various research directions as it combines tools and concepts related to the quantum-to-classical transition, quantum cosmology, quantum statistical mechanics and quantum stochastic processes.
@@ V.2 Outlook
For instance, while we argued that non-integrability could be a key factor, chaos has been also realized as detrimental for the emergence of classicality within the Wheeler-DeWitt equation due to the breakdown of the WKB approximaton and EID was invoked to remedy for that Calzetta and Gonzalez (1995); Cornish and Shellard (1998); Calzetta (2012). Within our non-relativistic quantum mechanical model we can not make any direct contribution to this question, but one view suggested by this work is that emergence of classicality is best viewed as a synergy of different mechanisms instead of a single all-ruling idea. Indeed, it has been recently pointed out that the emergence of classicality in the early Universe is still an unsolved puzzle Berjon et al. (2021), and the present perspective might be able to add a piece to it.
@@
Moreover, it would be desirable to get a better analytical understanding of the DHC. The results reported in Refs. Strasberg et al. (2023); Strasberg (2023) were restricted to three-time correlation functions and generalizing them to higher orders likely requires novel techniques (however, some progress for arbitrary $L$-time correlation functions in a different context was reported in Refs. Figueroa-Romero et al. (2019, 2021); Dowling et al. (2023a, b)). Further insights about the time scales at which decoherence arises are also desirable, and it is in particular also worth to ask about “recoherence times”: due to Poincaré recurrences the emergence of decoherence in a finite dimensional quantum system can not be a persistent, eternal phenomenon. In particular, it would be intriguing to find out how the recoherence time scales with the number of time steps $L$ or, more generally, with the net information aquired along a history. In addition, it seems that more research is necessary to understand the general consequences of approximate instead of exact decoherence given that the latter is not the rule.
@@
Finally, our model calculations presently suggest that for any coarse and slow observable of a non-integrable system every possible classical history is realized. While this suggests the realization of a wide array of histories, it is not necessarily true that “everything that can happen will happen” (as sometimes portrayed in both scientific and popular accounts of the MWI). For instance, extremely unlikely (sequences of) outcomes might be restricted to very small subspaces because they are highly atypical, but our results indicate that large subspaces are needed for the emergence of classicality, see also Ref. Strasberg and Schindler (2023). We also collected some evidence in Fig. 6 that there is structure in the Multiverse among different branches, but much more work in that direction needs to be done. Finally, whether all the perceived randomness in our world stems from quantum effects Albrecht and Phillips (2014) or not Del Santo and Gisin (2019) remains a fascinating question too.
Note added.—While this manuscript was under review, evidence appeared that confirms our results for realistic quantum-chaotic many-body systems and suggests that integrable finite-size systems show a much weaker form of decoherence Wang and Strasberg (2024).
@@ Acknowledgements
We are grateful to Anthony Aguirre, Josh Deutsch, Jonathan Halliwell, and Anastasiia Tiutiakina for discussions. PS is financially supported by “la Caixa” Foundation (ID 100010434, fellowship code LCF/BQ/PR21/11840014) and acknowledges further support from the European Commission QuantERA grant ExTRaQT (Spanish MICINN project PCI2022-132965), by the Spanish MINECO (project PID2019-107609GB-I00) with the support of FEDER funds, the Generalitat de Catalunya (project 2017-SGR-1127) and by the Spanish MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1). JS acknowledges support by MICIIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by the Generalitat de Catalunya.